Project

General

Profile

NDOS Commissioning Meeting 2011/08/29

Attendance: Jon, Peter, Martin, Enrique, Kurt, Mark Mathis, Mat, Minerba, Gary, Rafael, Susan, Tom, Andrew, Denis, Leon

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & NDOS Update

  • DSO scan this morning by Peter. Not clear if any additional
  • Reminder: power to NDSB will be turned off tomorrow morning at 9 am so that Chuck Andrews & Co. can install a monitor to investigate the grounding issue at NDOS; ndos-01 (computer in building) also needs to be shut down. Andrew and Denis will develop a procedure for the power outage tomorrow.
  • Mat will remove failed APDs immediately after the NDSB power outage tomorrow.
  • TDU: some confusion over the weekend as to which TDU we should be connect to. Will improve both documentation and make the start-up of the TDU client an icon.

DAQ Status & Plans

  • Plans for the week?
    • Peter: perhaps testing of the message analyzer (watchdog) application.
    • Kurt: will probably want to flatten the DAQ cfg. tables
  • APD Recovery (Leon):
    • able to test the seal between the board and spacer frame; last test to do of the ~20 APDs that look good so far. Assuming these are ok, then he'll ship these out to FNAL this week. The test takes ~10 minutes each.
    • found source of leak on the most recent failed APDs shipped to CalTech.

DCS Status & Plans

  • Martin: new alarm thresholds have been installed on CSS. Also, new overview is ready that will show "OR" of alarms. There is one FEB that has a temp. that oscillates between 0 and 40 deg. that is likely a hardware failure.
  • Water temp. will be raised to a degree or two to address an issue related to the minimum TECC drive current on some FEBs.

AOB

  • Denis: we should identify who will do the APD installation and start working on the procedures.